Going down a river, a boat went 12 km/h. Going up the river, the boat went only 2 km/h. How fast would the boat go if there was no current?
Found 2 solutions by mananth, Alan3354:
Answer by mananth(16949)

let speed of boat in still water be x
current speed be y
with curent speed = x+y = 12
against current speed = x-y =2
x+y=12
x-y=2
add the two equations
2x=14
x=7 speed of boat in still water

Answer by Alan3354(69443)

Going down a river, a boat went 12 km/h. Going up the river, the boat went only 2 km/h. How fast would the boat go if there was no current?
-------
The average (12 + 2)/2 = 7 km/hr